the length of a swimming pool is 8m longer than its width and the area is 105 meter square.

Answer by ewatrrr(24785)

 
Hi,
.
A = Lw
= (w+8)w = 105
w^2 + 8w -105 = 0 (Tossing out the negative solution for unit measure)
(w+15)(w-7) = 0, w = 7m and L = 15m